# Quantum computational speedup and retrocausality

### v3

@@ Abstract
Involving only the measurements of commuting observables – the problem-setting and the corresponding solution – quantum algorithms must be subject to classical logic. This allows us to flank their customary quantum description with a classical logic description. In the classical logic description of the quantum algorithm, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one (any one) of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. This is naturally a causal loop whose retrocausal character turns out to be implicit in the very notion of quantum state superposition, both an essential ingredient of the quantum computational speedup and one of the pillars of quantum mechanics. More in general, these causal loops turn out to be implicit in quantum correlations and, under the quantum cosmological assumption, also in the correlations of macroscopic reality.
1 Introduction
Let us first apologize for a longer-than-usual “Introduction”. It is to introduce a new way of viewing quantum computation that must be clarified from the outset.
1.1 Summary
@@
Note that this also establishes a very direct, quantitative, comparison between quantum and classical computation: in the optimal case, the number of computation steps needed by the quantum algorithm is that of a classical algorithm that benefits of the advanced knowledge of one of the possible halves of the information about the solution of the problem.
We called this the advanced knowledge rule, and showed that it fits all the major quantum algorithms $\left[4\right]$.
@@
In this work, besides providing an organic review of that approach, we should have demonstrated the advanced knowledge rule in question. We will show that not only it explains an optimal quantum speedup, it is also logically implied by it. This will be in the context of a novel approach to quantum algorithms. Involving only the measurements of commuting observables – the setting and the corresponding solution of the problem – they must be subject to classical logic. One can therefore flank their customary quantum description with a classical logic description. We will see that the latter highlights features that are implicit, but hidden, in their customary quantum description.
By the way, the need to time-symmetrize the ordinary quantum description of quantum algorithms will also be confirmed by a symmetry implicit in it. Since we are navigating in uncharted waters, this redundancy should not be despised.
Before going to show that the classical logic description of an optimal quantum speedup implies the advanced knowledge rule in question, let us make a few remarks on the latter.
@@
In today’s quantum information, the only way of knowing the number of computation steps (oracle queries) needed to solve an oracle problem in an optimal quantum way is finding the quantum algorithms that does that. Of course, this is an object of high ingenuity, uncertain achievement, and different from oracle problem to oracle problem. As far as we know, excluding extremely simple quantum algorithms with just one computation step, this object has been achieved only once, indeed in the case of the optimal quantum search algorithm of Grover $\left[6\right]$.
@@
Let us end this subsection by introducing the concept that stands at the basis of the present retrocausal explanation of the quantum speedup. It is a retrocausal explanation of the very notion of quantum state superposition, of course, both an essential ingredient of the quantum speedup (think of the notion of quantum parallel computation) an one of the pillars of quantum mechanics. Simply, its classical logic description necessarily involves retrocausality. Of course, it will turn out to be exactly the retrocausality involved in the present retrocausal explanation of the quantum speedup.
Let us resort to a pre-scientific notion of state superposition, which allows a very intuitive classical logic description of it. It is the legend of the monk seen by his disciples on the two banks of a river at the same time. Now, from a logical standpoint, these two possibilities, of being on one or the other bank, are mutually exclusive with one another. Apparently, being on both banks at the same time is a logical impossibility. However, instead, it becomes possible if one opens to retrocausality, as follows. The monk, first, follows a path that leads him to one of the two banks of the river; then (then in causal order) he goes sufficiently backward in time undoing the path in question (as in a reversed film), then he goes forward in time again but following a different path that brings him to the other bank. With this causality zigzag, he can, of course, have been seen on the two banks of the river at the same time.
We will see how the above classical logic description of a state superposition can be ported, as it is, into the quantum world where it explains the quantum speedup in a unified quantitative way for all oracle quantum computing, and more.
1.2 Classical logic description of Grover algorithm
For its special features, Grover’s quantum search algorithm will play a special role in the present work.
First, in its simplest instance, and any instance in Long’s version of it $\left[8\right]$, it provides the solution with certainty, what simplifies things.
Second, the solution is an invertible function of the setting of the problem – a condition for the reversibility of the quantum process in between: it will become important.
Third, it is an optimal quantum algorithm, namely one that demonstrably solves its problem with the minimum possible number of quantum computation steps $\left[9\right]$. This is obviously important when comparing the number of computation steps foreseen by the advanced knowledge rule in the optimal logical case with that taken by the quantum algorithm.
Fourth, it is a most universal quantum algorithm that can solve with quantum speedup any NP problem.
@@ 1.2 Classical logic description of Grover algorithm
By the way, the usual quantum description of Grover algorithm, as that of the other quantum algorithms, is limited to the process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is therefore (trivially) incomplete. The canonically complete quantum description of a quantum process is: initial measurement, unitary evolution, and final measurement. In view of what will follow, we complete it in the canonical way by adding to the usual process of computing and measuring the solution of the problem the upstream process of setting the problem.
Our “canonically complete” quantum description of the quantum algorithm is therefore as follows. An initial measurement of the problem-setting (e. g. the number of the drawer with the ball) in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random333 The problem-setter could unitarily change the problem-setting selected at random into any desired one. However, here, this would change nothing. . The unitary computation and measurement of the solution follow.
We will call this quantum description of the quantum algorithm, extended to the process of setting the problem, its ordinary quantum description, ordinary because it is still the causal quantum description of the evolutions of mainstream quantum mechanics.
By the way, further on we will have to further complete this “canonically complete quantum description” of the quantum algorithm, this time in a fundamental way.
@@
Then (“then” in causal order), she can produce the solution benefiting of that advanced knowledge. We should keep in mind that we are under a classical logic equivalent of what the quantum algorithm does. In it, the needed number of computation steps is that logically required. All this, of course, is in the mutually exclusive $or$ all the possible ways of knowing in advance half of the information about the solution. Eventually, note that this is exactly the advanced knowledge rule logically implied by the sheer existence of the present quantum speedup.
@@ 1.3 Completing the quantum description of Grover algorithm
Of course, there should be correspondence between the classical logic description of the quantum algorithm and its ordinary quantum description. However, its ordinary quantum description does not and, being causal in character, could not describe any causal loop. Coherently with the present line of thought, and despite the fact that it gives the correct number of computation steps needed to solve the problem (although in a different way for each quantum algorithm), we must conclude that it is incomplete. Of course, it should be completed in a way that satisfies the correspondence in question.
In view of what will follow, note that the process between the outcome of the initial measurement and the state immediately before the final measurement, being unitary, is in particular time-reversible. Since the final measurement outcome coincides with the state immediately before the final measurement (which gives the solution with certainty), we will speak, for short, of the unitary evolution between the initial and final measurement outcomes. Also note that this unitary evolution (i.e. the computation of the solution) leaves the eigenvalue selected by the initial measurement, i.e. the problem-setting, unaltered.
@@
As we will see, this time-symmetrization leaves the unitary part of the ordinary quantum description of the quantum algorithm (the computation of the solution) mathematically unaltered but changes the description of the behavior of causality along it. As seen by the problem-solver (we will explain this specification further below), the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. From now on, we will call the (physical) causal loops in question time-symmetric, because highlighted by a time-symmetrization procedure; we will call in the same way also the form of retrocausality entailed in them.
By the way, note that the retrocausality we have been dealing with is, in the first place, logical in character: it is the one present in the classical logic description of what the quantum algorithm does. Of course, retrocausality can have a purely logical form, all the paradoxes originated by it have this form; think of the grandfather paradox. However, with the time-symmetrization of the ordinary quantum description of the quantum algorithm, this logical retrocausality becomes physical, quantum (the quantum superposition of causal loops). Note that, obviously, this classical logic retrocausality would be physical also in the classical world; in the sense that also here to always find the solution by opening just one drawer would require knowing in advance half of the information about the solution produced in the future. However, of course, the possibility of having physical retrocausality in the classical world is another story. We will go back to this in the Discussion section. In the end, summing up, the logical form of retrocausality would always imply a corresponding physical form.
@@
Let us recall the classical logic description of the pre-scientific notion of state superposition: the monk being on the two banks of the river at the same time. Since the two possibilities are logically mutually exclusive with one another, this is apparently a logical impossibility. However, it is an impossibility that vanishes by resorting to retrocausality. After reaching one bank, the monk should sufficiently go backward in time by the inverse of the path followed to reach that bank and then forward in time again but on a different path that leads to the other bank.
Going to the quantum world, this kind of classical logic description of a state superposition can be applied, for example, to a photon passing in quantum superposition along the two branches of a Mach-Zehnder interferometer. The two branches should be tuned in such a way that the photon always ends up in the same detector. In this way, the whole process until “reaching” the detector (immediately before detection) is unitary and can therefore be run backward in time by the inverse transformation. Of course, there is a classical logic description of the photon (with its phase) reaching the detector having run either branch but these two possibilities are mutually exclusive with one another. However, there can be a classical logic description also in this case provided that one opens to retrocausality. After “reaching” the detector (immediately before the detection) having gone along one of the two branches, the photon should go backward in time, by the inverse of the time-forward unitary transformation, until before entering the Mack-Zehnder interferometer; then, going forward in time again, it should reach the same detector through the other branch. This time it can also be detected.
By the way, the above classical logic description of the behavior of the photon is ad hoc to rebuild its actual quantum behavior. The important thing is that this classical logic description is logically possible, which, of course, is.
Of course, the unitary quantum computation of the solution is a sequence of generations of quantum state superpositions and interferences. The former can be described in a classical logic way by resorting to the causality zigzags we are dealing with, the latter, of course, without any zigzag. After the simplifications, we should turn out with the advanced knowledge rule derived in a much more compact way by the classical logic description of the resulting quantum speedup. Now, about the advantage of relating the quantum speedup to retrocausality rather than quantum superposition, it is, of course, that of moving from a qualitative to a quantitative explanation of it.
@@ 1.4 Generalization
One is that the unitary evolution builds up one-to-one correlation between two initially unrelated (commuting) observables without changing the eigenvalue of the observable selected by the initial measurement. This building up can always occur with quantum speedup. This is, of course, the case of quantum algorithms but the unitary evolution between the two one-to-one correlated measurement outcomes may not be a quantum algorithm at all. In any case, that building up can always occur with (quadratic) quantum speedup. When we are not dealing with a quantum algorithm, obviously, the notion of problem solver and her advanced knowledge of half of the information about the solution vanishes. What instead survives is the reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place. It is this search that can consequently occur with quantum speedup.
@@
The other case is that there is no such building up because the two observables are already maximally entangled before the beginning of the quantum process. This is, for example, the case of quantum nonlocality. Consider the measurement (in the same basis) of two maximally entangled observables in a state of spatial separation between the two respective subsystems. Of course, there is one-to-one correlation between the two measurement outcomes and the unitary process of spatial separation does not change the eigenvalue selected by the initial measurement. As shown in $\left[5\right]$, time-symmetrizing the ordinary quantum description of quantum nonlocality (exactly as we did with that of quantum algorithms) yields a quantum superposition of causal loops that play the role of the hidden variables envisaged by Einstein and the others in their celebrated EPR paper $\left[10\right]$. Indeed, from the outcome of the first performed measurement (on one of the two subsystems) they locally go backward in time to the time the two subsystems were not yet space separate, where they locally change the state of the other subsystem in a way that ensures the appropriate correlation between the two future measurement outcomes. In other words, the causal loops we are dealing with exactly play the role of, or even be, the hidden variables envisaged by Einstein and the others.
@@
Note the many commonalities between the quantum speedup and quantum nonlocality that there are under the present theory: (i) both do something that is impossible to do in the classical case, respectively solving a problem with fewer computation steps than classically possible and instantaneously changing the state of an object at a distance, (ii) both do so by availing themselves of the present time-symmetric form of retrocausality, (iii) in both cases, the classical logic description of the quantum process highlights the incompleteness of its ordinary quantum description, (iv) in both cases, the ordinary quantum description is completed by time-symmetrizing it, and (v) in both cases we are dealing with a retrocausality internal to the quantum process that cannot be used by observers external to it to communicate with each other.

### v4

@@ Abstract
Involving only the measurements of commuting observables – the problem-setting and the corresponding solution – quantum algorithms should be subject to classical logic. This would allow flanking their customary quantum description with a classical logic description, with surprising consequences. In the classical logic description of the quantum algorithm, very simply, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. This is a causal loop whose retrocausal character turns out to be implicit in the very notion of quantum state superposition, both an essential ingredient of the quantum computational speedup and one of the pillars of quantum mechanics. Indeed, the key point of the work is that the classical logic description of a quantum state superposition must resort to a logical form of retrocausality that in turn must be physically implicit in the superposition itself. The existence of retrocausality in ordinary quantum physics implies e different way of viewing physical reality. It explains in a unified way all quantum speedups and quantum nonlocality. It highlights the teleological character of quantum algorithms, that is, their being evolutions toward a goal (the solution of the problem) with an attractor in the solution they will produce in the future (the solution again). Under the quantum cosmological assumption, it provides a plausible physical basis to the teleological character of natural evolutions.
1 Concept
The present work is little mathematics and mostly a different way of viewing the quantum computational speedup. Before going on to formalizations, we thought it appropriate to provide a conceptual exposition of the view.
1.1 Introduction
@@
We called this “the advanced knowledge rule”, and showed that it fits all the major quantum algorithms $\left[4\right]$. Note that it also establishes a very direct, quantitative, comparison between quantum and classical computation: in the optimal case, the number of computation steps needed by the quantum algorithm is that of a classical algorithm that benefits of the advanced knowledge of one of the possible halves of the information about the solution of the problem.
@@
By the way, the need to time-symmetrize the ordinary quantum description of quantum algorithms will also be confirmed by a symmetry implicit in this description. This goes in favor of the assumption that quantum algorithms are subject to classical logic and guarantees its consequence.
Until now, the advanced knowledge rule we are dealing with has been an empirical fact satisfied by all the quantum algorithms examined. In this work, we should have theoretically demonstrated it, incidentally in a surprisingly simple way. By the way, concerning retrocausality, it would be a demonstration with broad consequences on our way of viewing the physics of both the microscopic and the macroscopic phenomena.
We can start from the observation that, only involving the measurements of commuting observables – the problem-setting and the corresponding solution of the problem – quantum algorithms should be subject to classical logic. Therefore, their ordinary quantum description could be flanked by a classical logic description and, of course, the two descriptions should go along. As we will soon see, the classical logic description of what the quantum algorithm does – its producing a certain quantum speedup – very simply implies the advanced knowledge rule we are dealing with. Indeed, it immediately says that it is as if the the problem solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with a quantum speedup.
This is, of course, a causal loop, which in turn implies retrocausality. Naturally, it must be a retrocausality implicit – and hidden – in the ordinary quantum description of the quantum algorithm and made explicit by its classical logic description.
Now a quantum algorithm is a rather complex physical phenomenon. A natural question is then whether there is a more elementary quantum phenomenon whose classical logic description implies retrocausality. Of course, this is indeed the case. The phenomenon in question is quantum state superposition, both an essential ingredient of the quantum speedup and one of the pillars of quantum mechanics.
Let us show how its classical logic description must resort to retrocausality. This is shown in a most simple way by resorting to a prescientific notion of state superposition – it will be the same with the ordinary quantum description.
It is the ancient legend of the monk seen by his disciples on the two banks of a river at the same time. Now, from a logical standpoint, these two possibilities, of being on one or the other bank, are mutually exclusive with one another. Apparently, being on both banks at the same time is thus a logical impossibility. However, instead, it becomes possible if one opens to retrocausality, as follows. The monk, first, follows a path that leads him to one of the two banks of the river; then (“then” in causal order) he goes sufficiently backward in time undoing the path in question (as in a reversed film), then he goes forward in time again but following a different path that brings him to the other bank. With this causality zigzag, he can, of course, have been seen on the two banks of the river at the same time. We can also see that any classical logic description of this state superposition inevitably involves retrocausality. This is the rather simple key of the entire work.
We will see how the above classical logic description of a state superposition can be ported, as it is, into the quantum world where it explains the quantum speedup in a unified quantitative way for all oracle quantum computing, and much more.
1.2 Classical logic description of Grover algorithm
Before going to show that the classical logic description of an optimal quantum speedup implies the advanced knowledge rule we are dealing with, let us make a few remarks on the latter.
@@ 1.2 Classical logic description of Grover algorithm
In today’s quantum information, the only way of knowing the number of oracle queries needed to solve an oracle problem in an optimal quantum way is finding the quantum algorithms that does that. Of course, this is an object of high ingenuity, uncertain achievement, and different from oracle problem to oracle problem. As far as we know, excluding extremely simple quantum algorithms with just one computation step, this object has been achieved only once, indeed in the case of the optimal quantum search algorithm of Grover $\left[6\right]$.
@@
We go now to the derivation of the advanced knowledge rule by flanking the ordinary quantum description of Grover algorithm with its classical logic description.
By the way, for its special features, Grover’s quantum search algorithm will play a special role in the present work:
(i) In its simplest instance, and any instance in Long’s version of it $\left[8\right]$, it provides the solution with certainty, what simplifies things.
(ii) The solution is an invertible function of the setting of the problem – a condition for the reversibility of the quantum process in between: this will become important.
(iii) It is an optimal quantum algorithm, namely one that demonstrably solves its problem with the minimum possible number of quantum computation steps $\left[9\right]$. This is obviously important when comparing the number of computation steps foreseen by the advanced knowledge rule in the optimal logical case with that taken by the quantum algorithm.
(iv) It is a most universal quantum algorithm that can solve with quantum speedup any NP problem.
@@
By the way, we will adopt a quantum description of Grover algorithm different from the usual one, which, as that of the other quantum algorithms, is limited to the process of solving the problem (where the quantum speedup resides) and the final measurement of the solution. It is therefore (trivially) incomplete. The canonically complete quantum description of a quantum process is: initial measurement, unitary evolution, and final measurement. We therefore complete the quantum description of Grover algorithm by adding to the quantum description of the process of computing and measuring the solution of the problem that of the upstream process of setting the problem.
Our “canonically complete” quantum description of the quantum algorithm is therefore as follows. An initial measurement of the problem-setting (e. g. the number of the drawer with the ball) in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random333 The problem-setter could unitarily change the problem-setting selected at random into any desired one. However, here, this would change nothing. . The unitary computation and measurement of the solution follow. We will call this quantum description of the quantum algorithm, extended to the process of setting the problem, its ordinary quantum description, “ordinary” because it is still the causal quantum description of the evolutions of mainstream quantum mechanics.
By the way, further on we will have to further complete this “canonically complete” quantum description of the quantum algorithm, this time in a fundamental way.
@@
Then (“then” in causal order), she can produce the solution benefiting of that advanced knowledge. We should keep in mind that we are under a classical logic equivalent of what the quantum algorithm does. In it, the needed number of computation steps is that logically needed. All this, of course, is in the mutually exclusive $or$ all the possible ways of knowing in advance half of the information about the solution. Eventually, note that this is exactly the advanced knowledge rule logically implied by the sheer existence of the present quantum speedup.
@@ 1.3 Completing the quantum description of Grover algorithm
Of course, there should be correspondence between the classical logic description of the quantum algorithm and its ordinary quantum description. However, its ordinary quantum description does not and, being causal in character, could not describe any causal loop. Coherently with the present line of thought, and despite the fact that it gives the correct number of computation steps needed to solve the problem (although in an unfereseeably different way for each quantum algorithm), we must conclude that it is incomplete. Naturally, it should be completed in a way that satisfies the correspondence in question.
In view of what will follow, note that the process between the outcome of the initial measurement and the state immediately before the final measurement is both unitary and time-reversible. Since the final measurement outcome coincides with the state immediately before the final measurement (which gives the solution with certainty), we will also speak, for short, of the unitary evolution between the initial and final measurement outcomes, even if it comprises a final measurement that does not change the quantum state. Also note that this unitary evolution (i.e. the computation of the solution) leaves the eigenvalue selected by the initial measurement, i.e. the problem-setting, unaltered.
@@
As we will see, this time-symmetrization leaves the unitary part of the ordinary quantum description of the quantum algorithm (the computation of the solution) mathematically unaltered but changes the description of the behavior of causality along it. As seen by the problem-solver (we will explain this specification further below), the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. From now on, we will call the (physical) causal loops in question time-symmetric because highlighted by a time-symmetrization procedure; we will call in the same way also the form of retrocausality entailed in them.
By the way, note that the retrocausality we have been dealing with is, in the first place, logical in character: it is the one present in the classical logic description of what the quantum algorithm does. Of course, retrocausality can have a purely logical form, all the paradoxes originated by it have this form: think of the grandfather paradox. However, with the time-symmetrization of the ordinary quantum description of the quantum algorithm, this logical retrocausality becomes physical, quantum indeed (the quantum superposition of causal loops).
@@
Let us recall the classical logic description of the prescientific notion of state superposition: the monk being on the two banks of the river at the same time. Since the two possibilities are logically mutually exclusive with one another, this is apparently a logical impossibility. However, it is an impossibility that vanishes by resorting to retrocausality. After reaching one bank, the monk should sufficiently go backward in time by the inverse of the path followed to reach that bank and then forward in time again but on a different path that leads him to the other bank.
Going now to the quantum world, under the assumption that quantum algorithms can be described in a classical logic way, also the quantum superpositions that are an essential part of them should be describable in this way. Let us see how the above classical logic description of the monk being on the two banks of the river at the same time can be applied, for example, to a photon passing in quantum superposition along the two branches of a Mach-Zehnder interferometer.
The two branches should be tuned in such a way that the photon always ends up in the same detector. In this way, the whole process until “reaching” the detector (we mean: immediately before detection) is unitary and can therefore be run backward in time by the inverse transformation. Of course, there is a classical logic description of the photon (with its phase) “reaching” the detector having run either branch but these two possibilities are mutually exclusive with one another. However, there can be a classical logic description also in this case provided that one opens to retrocausality. After “reaching” the detector (immediately before detection) having gone along one of the two branches, the photon should go backward in time, by the inverse of the time-forward unitary transformation, until before entering the Mack-Zehnder interferometer; then, going forward in time again, it should reach the same detector through the other branch. This time it can also be detected.
This is a classical logic description of the photon going in quantum superposition through the upper and lower branch. One can see that any such description necessarily involves retrocausality.
Going now to the unitary quantum computation of the solution, this is a sequence of generations of quantum state superpositions and interferences. The former can thus be described in a classical logic way by resorting to the causality zigzags we are dealing with, the latter, of course, without any zigzag. After the simplifications, we should turn out with the advanced knowledge rule derived in a much more compact way by the classical logic description of the resulting quantum speedup. Now, about the advantage of relating the quantum speedup to retrocausality rather than quantum superposition, it is, of course, that of moving from a qualitative to a quantitative explanation of it.
@@ 1.4 Generalization
One is that the unitary evolution builds up one-to-one correlation between two initially unrelated (commuting) observables without changing the eigenvalue of the observable selected by the initial measurement. This building up can always occur with quantum speedup. This is, of course, the case of quantum algorithms but the unitary evolution between the two one-to-one correlated measurement outcomes may not be a quantum algorithm at all. In any case, that building up can always occur with (quadratic) quantum speedup. When we are not dealing with a quantum algorithm, obviously, the notion of problem solver and her advanced knowledge of half of the information about the solution vanishes. What instead survives is the reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place. It is this search what can occur with quantum speedup.
The other case is that there is no such building up because the two observables are already maximally entangled before the beginning of the quantum process. This is, for example, the case of quantum nonlocality.
@@
Consider the measurement (in the same basis) of two maximally entangled observables in a state of spatial separation between the two respective subsystems. Of course, there is one-to-one correlation between the two measurement outcomes and the unitary process of spatial separation does not change the eigenvalue selected by the initial measurement. As shown in $\left[5\right]$, time-symmetrizing the ordinary quantum description of quantum nonlocality (exactly as we did with that of quantum algorithms) yields a quantum superposition of causal loops that play the role of the hidden variables envisaged by Einstein and the others in their celebrated EPR paper $\left[10\right]$. Indeed, from the outcome of the first performed measurement (on one of the two subsystems) they locally go backward in time to the time the two subsystems were not yet space separate, where they locally change the state of the other subsystem in a way that ensures the appropriate correlation between the two future measurement outcomes. In other words, the causal loops we are dealing with exactly play the role of, or even be, the hidden variables envisaged by Einstein and the others.
@@
Note the many commonalities between the quantum speedup and quantum nonlocality that there are under the present theory, given that both do something that is impossible to do in the classical case, respectively solving a problem with fewer computation steps than classically possible and instantaneously changing the state of an object at a distance:
(i) both do so by availing themselves of the present time-symmetric form of retrocausality,
(ii) in both cases, the classical logic description of the quantum process highlights the incompleteness of its ordinary quantum description,
(iii) in both cases, the ordinary quantum description is completed by time-symmetrizing it, and
(iv) in both cases we are dealing with a retrocausality internal to the quantum process that cannot be used by observers external to it, for example to communicate backward in time or instantaneously at a distance.
@@
The reader not interested in further technical detail or the formalizations of what exposed until now can go to Subsection 4.2. “Looking forward”.